Write thd equation of the line passing through (4,7) and parallel to 6x-7y=10
Answer by ewatrrr(24785)

 
Hi,
equation of the line passing through (4,7) and parallel to 6x-7y=10 OR y = (6/7)x -10/7
y = (6/7)x + b
7 = (6/7)4 + b ||Using pt(4,7) to solve for b
25/7 = b
y = (6/7)x + 25/7 or in Standard form: 6x-7y = -25